OpenStudy (yanasidlinskiy):

Derive v^2/r from Ohm's law

OpenStudy (agent0smith):

You just need V = IR and P = IV First, solve V = IR for I - just divide both sides by R\[\large I = \frac{ V }{ R }\] Now replace the I in P = IV with that \[\large P = \frac{ V }{ R }*V\]and that's it.

OpenStudy (yanasidlinskiy):

so it would be p=v/r*v?

OpenStudy (agent0smith):

Well you simplify it... V times V = ?

OpenStudy (yanasidlinskiy):

V^2

OpenStudy (yanasidlinskiy):

so v^s/R?

OpenStudy (yanasidlinskiy):

*^2

OpenStudy (agent0smith):

Well yes, since that is what you were deriving... :P

OpenStudy (yanasidlinskiy):

So the final is... p=v/r*v

OpenStudy (yanasidlinskiy):

?

OpenStudy (agent0smith):

no lol you just wrote the final expression!

OpenStudy (yanasidlinskiy):

LOL

OpenStudy (yanasidlinskiy):

I don't get it. That's so weird. I'll never understand physics

OpenStudy (agent0smith):

It's just algebra! V*V/R is V^2/R!

OpenStudy (yanasidlinskiy):

Ik that. so for it being derived it would be v*v/R Right??

OpenStudy (agent0smith):

NO GODDAMMIT

OpenStudy (agent0smith):

The derivation is the whole thing, the first post!

OpenStudy (yanasidlinskiy):

WTF! I"M FCKN CONFUSED DUMAS

OpenStudy (agent0smith):

In my first post, the derivation is basically the whole thing, at least from the word "First,"